OpenStudy (anonymous):

A glider with mass m = 0.210kg sits on a frictionless horizontal air track, connected to a spring with force constant k = 5.20N/m . You pull on the glider, stretching the spring 0.130m , and then release it with no initial velocity. The glider begins to move back toward its equilibrium position (x=0).

OpenStudy (anonymous):

Is there more?

OpenStudy (anonymous):

What is the speed of the glider when it returns to x=0?

OpenStudy (anonymous):

What must the initial displacement of the glider be if its maximum speed in the subsequent motion is to be 3.00m/s ?

OpenStudy (anonymous):

what's the energy in the spring when it's stretched .130 m?

OpenStudy (anonymous):

any ideas?

OpenStudy (anonymous):

nope. :(

OpenStudy (anonymous):

do you have any solution that can guide me?

OpenStudy (anonymous):

are you doing work/energy theorem?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

cool.

OpenStudy (anonymous):

i got 3 zero's already . -_- i cannot have any mistake now. it will fail me -_-

OpenStudy (anonymous):

you should be given an expression for the energy a spring stores when compressed or stretched if not you can integrate to find the work required to compress/stretch \[W =\int\limits_{ }^{ } F*dx = \int\limits_{ }^{} -kx*dx = ?\]

OpenStudy (anonymous):

Any luck finding the expression in your notes or text? Or, alternately, evaluating the integral?

OpenStudy (anonymous):

thats the given only. its an online physics assignment so its so hard -_- the given is only that. oh my god -_-

OpenStudy (anonymous):

University of Phoenix?

OpenStudy (anonymous):

sears and zemansky's university of physics with modern physics

OpenStudy (anonymous):

thanks:)

OpenStudy (anonymous):

Did you find the expression for energy in a spring?

OpenStudy (anonymous):

not yet. this subject is so difficult. and integral also -_- im doom.